# Exact Maximin Share Fairness via Adjusted Supply

### v1

@@ Abstract
We prove that, for division of $m$ goods among $n$ agents with monotone valuations, there always exists an assignment of subsets of goods to the agents such that they receive at least their maximin shares and no single good is allocated to more than $3\log m$ agents. In addition, the sum of the sizes of the assigned subsets—i.e., the total number of goods assigned, with copies—does not exceed $m$. Further, for additive valuations, we prove that there always exists an MMS assignment in which no single good is allocated to more than $2$ agents and the total number of goods assigned, with copies, is at most $2m$. For additive ordered valuations, we obtain an upper bound of $O(\sqrt{\log m})$ on the maximum assignment multiplicity across goods and an $m+\widetilde{O}\left(\frac{m}{\sqrt{n}}\right)$ bound for the total number of goods assigned.
For fair division of chores, we upper bound the number of chores that need to be discarded for ensuring MMS fairness. We prove that, under monotone costs, there always exists an MMS assignment in which, out of the $m$ given chores, at most $\frac{m}{e}$ remain unassigned. We further show that, under additive costs, one can ensure MMS fairness by discarding at most $\frac{2m}{11}+n$ chores. For additive ordered costs, we establish that MMS fairness can be achieved while keeping at most $\widetilde{O}\left(\frac{m}{n^{1/4}}\right)$ chores unassigned. We also prove that the obtained bounds for monotone valuations and monotone costs are essentially tight.
@@ 1 Introduction
Additive Ordered Valuations
(Theorem 3.4)
$12\sqrt{\log m}$
$m+\frac{6m\sqrt{\log m}}{\sqrt{n}}$
@@
  • This work also obtains specialized bounds for instances in which the agents’ valuations are additive and all the agents rank (with respect to their individual valuations) the $m$ goods in the same order. For additive ordered valuations, we establish an $\ell_{\infty}$ bound of $O(\sqrt{\log m})$ and an $\ell_{1}$ bound of $m+\widetilde{O}\left(\frac{m}{\sqrt{n}}\right)$; where, $\widetilde{O}(\cdot)$ subsumes logarithmic dependencies. Note that here the $\ell_{1}$ norm additively exceeds the given number of goods, $m$, by only a lower-order term.
@@
Additive Ordered Costs
(Theorem 4.4)
$\frac{3m\left(\log m\right)^{\nicefrac{{3}}{{2}}}}{n^{\nicefrac{{1}}{{4}}}}$
@@
  • We further show that, under additive costs, one can ensure MMS fairness by discarding at most $\frac{2m}{11}+n$ chores.
  • For additive ordered costs, we establish that MMS fairness can be achieved while keeping $\widetilde{O}\left(\frac{m}{n^{1/4}}\right)$ chores unassigned.
@@ 2 Notation and Preliminaries
A fair division instance is given by a tuple $\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, where $[n]=\{1,2,.\dots,n\}$ is the set of $n\in\mathbb{Z}_{+}$ agents, $[m]=\{1,2,\dots,m\}$ the set of $m\in\mathbb{Z}_{+}$ indivisible goods, and for each agent $i\in[n]$, the set function $v_{i}:2^{[m]}\to\mathbb{R}_{+}$ denotes the valuation of agent $i$ over subsets of goods. Specifically, $v_{i}(S)\in\mathbb{R}_{+}$ denotes the value that agent $i$ derives from the subset $S\subseteq[m]$ of goods.
@@
We will additionally consider instances wherein all the agents’ valuations are additive and they rank the goods in the same order. That is, there exists an indexing of the $m$ goods $g_{1},g_{2},\ldots,g_{m}$ with the property that $v_{i}(g_{1})\geq v_{i}(g_{2})\geq\ldots\geq v_{i}(g_{m})$ for all the agents $i\in[n]$. We will refer to such settings as instances with additive ordered valuations.
@@ 3.1 Monotone Valuations
3.2 Additive Valuations
@@
Proof.
3.3 Additive Ordered Valuations
This section addresses additive valuations that rank the goods in the same order. Thoughout this section, we will write $g_{1},g_{2},\ldots,g_{m}$ to denote the indexing of the goods that satifies $v_{i}(g_{1})\geq v_{2}(g_{2})\geq\ldots\geq v_{i}(g_{m})$ for all agents $i\in[n]$. Theorem 3.4 (stated below) provides our main result for additive ordered valuations.
Remark.
It is relevant to note that prior works on MMS, for additive valuations, assume that one can restrict attention to ordered instances without loss of generality. This follows from a reduction of Bouveret and Lemaître [7] via which, and for any instance $\mathcal{I}$ with additive valuations, one can construct an ordered instance $\mathcal{I}^{\prime}$ such that if $\mathcal{I}^{\prime}$ admits an MMS allocation $\mathcal{A}^{\prime}$ then, in fact, $\mathcal{I}$ also admits an MMS allocation $\mathcal{A}$ [^6] See [6] for an approximation-preserving version of this result. The reduction works by deriving from $\mathcal{A}^{\prime}$ a picking sequence over the agents and then showing that greedily assigning goods in $\mathcal{I}$, via this sequence, leads to an MMS allocation $\mathcal{A}$.
At the same time, we note that Theorem 3.4 generalizes if one were to relinquish the requirement that each agent $i$’s bundle, $A_{i}$, must be a subset of $[m]$. That is, in contrast to the rest of the paper, if $A_{i}$ is allowed to be a multiset, in which each extra copy of any good $g$ fetches an additional value of $v_{i}(g)$, then the reduction from [7] applies and the guarantee given in Theorem 3.4 extends to all additive valuations.
Theorem 3.4.
Every fair division instance $\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$ with additive ordered valuations admits an MMS multi-allocation $\mathcal{A}=(A_{1},\ldots A_{n})$ in which no single good is assigned to more than $12\sqrt{\log m}$ agents and the total number of goods assigned, with copies, is at most $m+\frac{6m\sqrt{\log m}}{\sqrt{n}}$. That is, the characteristic vector $\chi^{\mathcal{A}}$ of $\mathcal{A}$ satisfies
@@ 3.3 Additive Ordered Valuations
Recall that in the case of ordered valuations, we have an indexing $g_{1},g_{2},\ldots,g_{m}$ over the goods under which $v_{i}(g_{1})\geq v_{i}(g_{2})\geq\ldots\geq v_{i}(g_{m})$ for all the agents $i\in[n]$.
We will now define a probabilistic event $G_{2}$ considering the dyadic prefixes of $\{g_{1},g_{2},\ldots,g_{m}\}$ and the random multi-allocation $\mathcal{R}$. For each $t\in\{0,1,\ldots,\log m\}$, define the prefix $P_{t}\coloneqq\{g_{1},g_{2},\ldots,g_{2^{t}}\}$ to be the set of the $2^{t}$ most valuable goods. In addition, let $\mathcal{P}$ be the collection of all such prefixes, $\mathcal{P}\coloneqq\{P_{0},P_{1},\dots,P_{\log m}\}$.
@@
Therefore, for each dyadic prefix subset $P\in\mathcal{P}$ it holds that
@@
Recall the goods are indexed in decreasing order of value, $v_{i}(g_{1})\geq v_{i}(g_{2})\ldots\geq v_{i}(g_{m})$. Also, the collection $\mathcal{P}=\{P_{0},P_{1},\dots,P_{\log m}\}$ contains all the dyadic prefixes, $P_{t}=\{g_{1},g_{2},\ldots,g_{2^{t}}\}$ with $t\in\{0,1,\dots,\log m\}$. We will construct the desired multi-allocation $\mathcal{A}$ from $\mathcal{B}$ by systematically replacing copies of a lower-valued good, $\ell$, by those of a higher-valued good $h$; see Algorithm 1 for details.
@@
In any fair division instance with additive ordered valuations, let $\mathcal{B}$ be an MMS multi-allocation that satisfies the properties stated in Corollary 3.7. Then, from $\mathcal{B}$, we can construct an MMS multi-allocation $\mathcal{A}$ in the given instance such that $\left\|\chi^{\mathcal{A}}\right\|_{\infty}\leq 12\sqrt{\log m}$ and $\left\|\chi^{\mathcal{A}}\right\|_{1}=\left\|\chi^{\mathcal{B}}\right\|_{1}$.
@@
3:     Let $\ell\in[m]$ be the lowest-indexed good with $\chi^{\mathcal{A}}_{{\ell}}>12\sqrt{\log m}$.
4:     Let $h\in[m]$ be any good of lower index (equivalently, higher value) than $\ell$ with $\chi^{\mathcal{A}}_{h}<12\sqrt{\log m}$. {We show that such a good $h$ always exists.}
5:     Pick an agent $i$ such that good $\ell\in A_{i}$ and $h\notin A_{i}$. {Since $\chi^{\mathcal{A}}_{h}<\chi^{\mathcal{A}}_{{\ell}}$, such an agent $i$ necessarily exists.}
26:     Update $A_{i}\leftarrow\left(A_{i}\setminus\{\ell\}\right)\cup\{h\}$.
{Each iteration of the while loop reduces the number of copies of a lower-valued good $\ell$ by commensurately increasing the number of copies of a higher-valued good $h$.}
@@
Algorithm 1 Constructing $\mathcal{A}$ from $\mathcal{B}$
@@
We will first show that the while-loop in Algorithm 1 successfully executes for each maintained multi-allocation $\mathcal{A}$ with $\left\|\chi^{\mathcal{A}}\right\|_{\infty}>12\sqrt{\log m}$. In particular, we will prove that if $\left\|\chi^{\mathcal{A}}\right\|_{\infty}>12\sqrt{\log m}$, for a maintained $\mathcal{A}$, then the good $h$ sought in Line 4 necessarily exists. Also, note that the while-loop must terminate after polynomially many iterations, since the following potential strictly decreases in each iteration: $\sum_{g\in[m]}\ \max\{0,\chi^{\mathcal{A}}_{g}-12\sqrt{\log m}\}$. Hence, for the returned multi-allocation $\mathcal{A}$ (obtained after the while-loop terminates), we have $\left\|\chi^{\mathcal{A}}\right\|_{\infty}\leq 12\sqrt{\log m}$.
Consider any iteration of the while-loop. Let $\mathcal{A}$ be the current multi-allocation (with $\left\|\chi^{\mathcal{A}}\right\|_{\infty}>12\sqrt{\log m}$) and $\ell$ be the good selected in Line 3. Assume here, towards a contradiction, that the good $h$ desired in Line 4 does not exist. That is, for each good $k\in L_{\ell}$, we have $\chi^{\mathcal{A}}_{k}\geq 12\sqrt{\log m}$; here, $L_{\ell}$ denotes the subset of all the goods with index at most that of $\ell$. By definition, $\ell\in L_{\ell}$. Also, note that, the selection criterion in Line 3 ensures that all the goods’ reassignments that have happened before the current iteration must have been between goods in $L_{\ell}$. That is, each $\ell^{\prime}$ and $h^{\prime}$ considered in any previous iteration satisfy $\ell^{\prime},h^{\prime}\in L_{\ell}$. This observation implies that, in current multi-allocation $\mathcal{A}$ and the initial multi-allocation $\mathcal{B}$, the total number of assignments (with copies) of the goods in $L_{\ell}$ are equal:
$\sum_{k\in L_{\ell}}\chi^{\mathcal{A}}_{k}=\sum_{k\in L_{\ell}}\chi^{\mathcal{B}}_{k}$ (12)
@@
Since $L_{\ell}\subseteq P_{t+1}$, inequality (13) implies $\sum_{k\in P_{t+1}}\chi^{\mathcal{B}}_{k}>6\sqrt{\log m}|P_{t+1}|$. This bound, however, contradicts the fact that multi-allocation $\mathcal{B}$ upholds property (ii) in Corollary 3.7 for $P_{t+1}\in\mathcal{P}$.
@@
Further, note that, in each iteration of the while-loop, the cardinality of the bundle $A_{i}$ is maintained – we add a higher-valued good, $h$, to it and remove a lower-valued one, $\ell$. Hence, for each agent $i$, the following inequalities continue to hold throughout the execution of the algorithm: $v_{i}(A_{i})\geq v_{i}(B_{i})$ and $|A_{i}|=|B_{i}|$. Therefore, for the returned multi-allocation $\mathcal{A}$, the stated $\ell_{1}$ bound holds: $\left\|\chi^{\mathcal{A}}\right\|_{1}=\sum_{i=1}^{n}|A_{i}|=\sum_{i=1}^{n}|B_{i}|=\left\|\chi^{\mathcal{B}}\right\|_{1}$. Finally, the facts that $\mathcal{B}$ is an MMS multi-allocation and $v_{i}(A_{i})\geq v_{i}(B_{i})$, for each $i$, imply that $\mathcal{A}$ is also MMS. Overall, we have that the returned allocation $\mathcal{A}$ is MMS, and it satisfies the stated $\ell_{1}$ and $\ell_{\infty}$ bounds. The lemma stands proved. ∎

### v2

@@ Abstract
We prove that, for division of $m$ goods among $n$ agents with monotone valuations, there always exists an assignment of subsets of goods to the agents such that they receive at least their maximin shares and no single good is allocated to more than $3\log m$ agents. In addition, the sum of the sizes of the assigned subsets—i.e., the total number of goods assigned, with copies—does not exceed $m$. We also address instances with identically ordered valuations. In such instances, the marginal values of the goods conform to a common order across all the agents and the subsets. For identically ordered valuations, we obtain an upper bound of $O(\sqrt{\log m})$ on the maximum assignment multiplicity across goods and an $m+\widetilde{O}\left(\frac{m}{\sqrt{n}}\right)$ bound for the total number of goods assigned. Further, for additive valuations, we prove that there always exists an MMS assignment in which no single good is allocated to more than $2$ agents and the total number of goods assigned, with copies, is at most $2m$.
For fair division of chores, we upper bound the number of chores that need to be discarded for ensuring MMS fairness. We prove that, under monotone costs, there always exists an MMS assignment in which, out of the $m$ given chores, at most $\frac{m}{e}$ remain unassigned. For identically ordered costs, we establish that MMS fairness can be achieved while keeping at most $\widetilde{O}\left(\frac{m}{n^{1/4}}\right)$ chores unassigned. We further show that, under additive costs, one can ensure MMS fairness by discarding at most $\frac{2m}{11}+n$ chores. We also prove that the obtained bounds for monotone valuations and monotone costs are essentially tight.
@@ 1 Introduction
Identically Ordered Valuations
(Theorem 3.4)
$12\sqrt{\log m}$
$m+\frac{6m\sqrt{\log m}}{\sqrt{n}}$
@@
  • This work also obtains specialized bounds for instances in which the marginal values of the goods conform to a common order across all the agents and the subsets. In particular, we say that the valuations are identically ordered if there exists an indexing of the $m$ goods, $\{g_{1},\ldots g_{m}\}$, such that for each pair of goods $g_{s},g_{t}\in[m]$, with index $s<t$, and any $S\subset[m]$ that does not contain $g_{s}$ and $g_{t}$, each agent values $S\cup\{g_{s}\}$ at least as much as $S\cup\{g_{t}\}$. For such identically ordered valuations, we establish an $\ell_{\infty}$ bound of $O(\sqrt{\log m})$ and an $\ell_{1}$ bound of $m+\widetilde{O}\left(\frac{m}{\sqrt{n}}\right)$; where, $\widetilde{O}(\cdot)$ subsumes logarithmic dependencies. Note that here the $\ell_{1}$ norm additively exceeds the given number of goods, $m$, by only a lower-order term.
@@
Identically Ordered Costs
(Theorem 4.4)
$\frac{3m\left(\log m\right)^{\nicefrac{{3}}{{2}}}}{n^{\nicefrac{{1}}{{4}}}}$
@@
  • For identically ordered costs, we establish that MMS fairness can be achieved while keeping $\widetilde{O}\left(\frac{m}{n^{1/4}}\right)$ chores unassigned.
  • We further show that, under additive costs, one can ensure MMS fairness by discarding at most $\frac{2m}{11}+n$ chores
@@ 2 Notation and Preliminaries
A fair division instance is given by a tuple $\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, where $[n]=\{1,2,.\dots,n\}$ is the set of $n\in\mathbb{Z}_{+}$ agents, $[m]=\{1,2,\dots,m\}$ the set of $m\in\mathbb{Z}_{+}$ indivisible goods, and for each agent $i\in[n]$, the set function $v_{i}:2^{[m]}\to\mathbb{R}_{+}$ denotes the valuation of agent $i$ over subsets of goods. Specifically, $v_{i}(S)\in\mathbb{R}_{+}$ denotes the value that agent $i$ derives from the subset $S\subseteq[m]$ of goods. For subsets $S\subseteq[m]$ and $g\in[m]$, we will write $S+g$ to denote the union $S\cup\{g\}$.
@@
We will additionally consider instances with identically ordered valuations. Here, we have an indexing of the $m$ goods, $\{g_{1},\ldots g_{m}\}$, such that for each pair of goods $g_{s},g_{t}$, with index $s<t$, and all agents $i\in[n]$, the inequality $v_{i}(S+g_{s})\geq v_{i}(S+g_{t})$ holds for each subset $S\subset[m]$ that does not contain $g_{s}$ and $g_{t}$; see Example 3.3 in Section 3.2.
@@ 3.1 Monotone Valuations
3.2 Identically Ordered Valuations
Recall that valuations in a fair division instance are said to be identically ordered if there exists an indexing $\{g_{1},\ldots g_{m}\}$ of the goods such that for each pair of goods $g_{s},g_{t}$, with index $s<t$, and all agents $i\in[n]$, the inequality $v_{i}(S+{g_{s}})\geq v_{i}(S+{g_{t}})$ holds for each subset $S\subset[m]$ that does not contain $g_{s}$ and $g_{t}$. Note that additive ordered valuations are identically ordered. The following example highlights that identically ordered valuations are not confined to be subadditive, or even superadditive.
Example 3.3.
Let $g_{1},\dots,g_{m}$ be a fixed indexing of the set $[m]$ of goods, and, for each agent $i$, let $w_{i}:[m]\to\mathbb{R}_{+}$ be weights on the goods that satisfy $w_{i}(g_{1})\geq w_{i}(g_{2})\geq\dots\geq w_{i}(g_{m})$. Note that, for all the agents, the weights respect the common indexing. For each $i\in[n]$, let $f_{i}:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ be a monotone nondecreasing function. Define valuation $v_{i}(S)\coloneqq f_{i}\left(\sum_{g\in S}w_{i}(g)\right)$, for each subset $S\subseteq[m]$ and agent $i$. Since the functions $f_{i}$s are monotone nondecreasing, the valuations $v_{i}$s are identically ordered: for all agents $i$, the following inequality holds for each pair of goods $g_{s},g_{t}$, with index $s<t$, and each subset $S$ that does not contain $g_{s}$ and $g_{t}$:
$v_{i}(S+g_{s})=f_{i}\left(\sum_{g\in S}w_{i}(g)\ +\ w_{i}(g_{s})\right)\geq f_{i}\left(\sum_{g\in S}w_{i}(g)\ +\ w_{i}(g_{t})\right)=v_{i}(S+g_{t}).$
With $f_{i}$s as the identify function, we obtain additive ordered valuations. Setting $f_{i}(w)=\sqrt{w}$ gives us subadditive valuations. On the other hand, with $f_{i}(w)=\exp(w)$, we obtain superadditive valuations. In all of these cases, our result for identically ordered valuations holds.
Theorem 3.4 (stated below) provides our upper bounds on the supply adjustments for MMS under identically ordered valuations.
@@ 3.2 Identically Ordered Valuations
Every fair division instance $\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$ with identically ordered valuations admits an MMS multi-allocation $\mathcal{A}=(A_{1},\ldots A_{n})$ in which no single good is assigned to more than $12\sqrt{\log m}$ agents and the total number of goods assigned, with copies, is at most $m+\frac{6m\sqrt{\log m}}{\sqrt{n}}$. That is, the characteristic vector $\chi^{\mathcal{A}}$ of $\mathcal{A}$ satisfies
@@
Let $g_{1},\dots,g_{m}$ be an indexing of the set $[m]$ of goods that satisfies to property laid out in LABEL:def:identically-ordered. We will now define a probabilistic event $G_{2}$ considering the dyadic prefixes of $\{g_{1},g_{2},\ldots,g_{m}\}$ and the random multi-allocation $\mathcal{R}$. For each $t\in\{0,1,\ldots,\log m\}$, define the prefix $P_{t}\coloneqq\{g_{1},g_{2},\ldots,g_{2^{t}}\}$ to be the set of the $2^{t}$ lowest-indexed goods (recall that these have the highest marginal values). In addition, let $\mathcal{P}$ be the collection of all such prefixes, $\mathcal{P}\coloneqq\{P_{0},P_{1},\dots,P_{\log m}\}$.
@@
Therefore, for each dyadic prefix subset $P\in\mathcal{P}$, it holds that
@@
Recall that, under identically ordered valuations, the goods are indexed $g_{1},\ldots,g_{m}$ in decreasing order of marginal values. That is, for each pair of goods $g_{s}$ and $g_{t}$, with index $s<t$, we have $v_{i}(S+g_{s})\geq v_{i}(S+g_{t})$ for every agent $i$ and each subset $S$ that does not contain $g_{s}$ and $g_{t}$. Also, the collection $\mathcal{P}=\{P_{0},P_{1},\dots,P_{\log m}\}$ contains all the dyadic prefixes, $P_{t}=\{g_{1},g_{2},\ldots,g_{2^{t}}\}$ with $t\in\{0,1,\dots,\log m\}$. We will construct the desired multi-allocation $\mathcal{A}$ from $\mathcal{B}$ by systematically replacing copies of a good $\ell$ with lower marginal value, by those of a good $h$ with higher marginal value; see Algorithm 1 for details.
@@
In any fair division instance with identically ordered valuations, let $\mathcal{B}$ be an MMS multi-allocation that satisfies the properties stated in Corollary 3.7. Then, from $\mathcal{B}$, we can construct an MMS multi-allocation $\mathcal{A}$ in the given instance such that $\left\|\chi^{\mathcal{A}}\right\|_{\infty}\leq 12\sqrt{\log m}$ and $\left\|\chi^{\mathcal{A}}\right\|_{1}=\left\|\chi^{\mathcal{B}}\right\|_{1}$.
@@
3:     Let $g_{t}\in[m]$ be the good with the lowest index $t$ that satisfies $\chi^{\mathcal{A}}_{g_{t}}>12\sqrt{\log m}$.
4:     Let $g_{s}\in[m]$ be any good of lower index $s<t$ (equivalently, higher marginal value) with $\chi^{\mathcal{A}}_{g_{s}}<12\sqrt{\log m}$. {We show that such a good $g_{s}$ always exists.}
5:     Pick an agent $i$ such that good $g_{t}\in A_{i}$ and $g_{s}\notin A_{i}$. {Since $\chi^{\mathcal{A}}_{g_{s}}<\chi^{\mathcal{A}}_{{g_{t}}}$, such an agent $i$ necessarily exists.}
26:     Update $A_{i}\leftarrow\left(A_{i}\setminus\{g_{t}\}\right)\cup\{g_{s}\}$.
{This iteration of the while loop reduces the number of copies of good $g_{t}$ by commensurately increasing the number of copies of good $g_{s}$, where $s<t$.}
@@
Algorithm 1 Copy Redistribution for Identically Ordered Valuations
@@
We will first show that the while-loop in Algorithm 1 successfully executes for each maintained multi-allocation $\mathcal{A}$ with $\left\|\chi^{\mathcal{A}}\right\|_{\infty}>12\sqrt{\log m}$. In particular, we will prove that if $\left\|\chi^{\mathcal{A}}\right\|_{\infty}>12\sqrt{\log m}$, for a maintained $\mathcal{A}$, then the good $g_{s}$ sought in Line 4 necessarily exists. Also, note that the while-loop must terminate after polynomially many iterations, since the following potential strictly decreases in each iteration: $\sum_{g\in[m]}\ \max\{0,\chi^{\mathcal{A}}_{g}-12\sqrt{\log m}\}$. Hence, for the returned multi-allocation $\mathcal{A}$ (obtained after the while-loop terminates), we have $\left\|\chi^{\mathcal{A}}\right\|_{\infty}\leq 12\sqrt{\log m}$.
Consider any iteration of the while-loop. Let $\mathcal{A}$ be the current multi-allocation (with $\left\|\chi^{\mathcal{A}}\right\|_{\infty}>12\sqrt{\log m}$) and $g_{t}$ be the good selected in Line 3. Assume here, towards a contradiction, that the good $g_{s}$ desired in Line 4 does not exist. That is, for each good $k\in L_{t}$, we have $\chi^{\mathcal{A}}_{k}\geq 12\sqrt{\log m}$; here, $L_{t}$ denotes the subset of all the goods with index at most $t$. By definition, $g_{t}\in L_{t}$. Also, note that, the selection criterion in Line 3 ensures that all the goods’ reassignments that have happened before the current iteration must have been between goods in $L_{t}$. That is, each pair of goods $g_{t^{\prime}}$ and $g_{s^{\prime}}$ considered in any previous iteration satisfy $g_{t^{\prime}},g_{s^{\prime}}\in L_{t}$. This observation implies that, in the current multi-allocation $\mathcal{A}$ and the initial multi-allocation $\mathcal{B}$, the total number of assignments (with copies) of the goods in $L_{t}$ are equal:
$\sum_{k\in L_{t}}\chi^{\mathcal{A}}_{k}=\sum_{k\in L_{t}}\chi^{\mathcal{B}}_{k}$ (11)
@@
Since $L_{t}\subseteq P_{\ell+1}$, inequality (12) implies $\sum_{k\in P_{\ell+1}}\chi^{\mathcal{B}}_{k}>6\sqrt{\log m}|P_{\ell+1}|$. This bound, however, contradicts the fact that multi-allocation $\mathcal{B}$ upholds property (ii) in Corollary 3.7 for $P_{\ell+1}\in\mathcal{P}$.
@@
Further, note that, in each iteration of the while loop, the cardinality of the bundle $A_{i}$ is maintained – we add a lower-indexed good, $g_{s}$, to it and remove a higher-indexed one, $g_{t}$. Since the valuations are identically ordered, in any such replacement, the updated value of agent $i$’s bundle, $v_{i}((A_{i}\setminus\{g_{t}\})\cup\{g_{s}\})$, is at least its original value $v_{i}(A_{i})=v_{i}((A_{i}\setminus\{g_{t}\})\cup\{g_{t}\})$. Hence, for each agent $i$, the following inequalities continue to hold: $v_{i}(A_{i})\geq v_{i}(B_{i})$ and $|A_{i}|=|B_{i}|$. Therefore, for the returned multi-allocation $\mathcal{A}$, the stated $\ell_{1}$ bound holds: $\left\|\chi^{\mathcal{A}}\right\|_{1}=\sum_{i=1}^{n}|A_{i}|=\sum_{i=1}^{n}|B_{i}|=\left\|\chi^{\mathcal{B}}\right\|_{1}$. Finally, the facts that $\mathcal{B}$ is an MMS multi-allocation and $v_{i}(A_{i})\geq v_{i}(B_{i})$, for each $i$, imply that $\mathcal{A}$ is also MMS. Overall, we have that the returned allocation $\mathcal{A}$ is MMS, and it satisfies the stated $\ell_{1}$ and $\ell_{\infty}$ bounds. The lemma stands proved. ∎
@@
Remark.
It is relevant to note that prior works on MMS, for additive valuations, assume that one can restrict attention, without loss of generality, to instances with valuations that identically ordered in addition to being additive. This follows from a reduction of Bouveret and Lemaître [7] via which, and for any instance $\mathcal{I}$ with additive valuations, one can construct an instance $\mathcal{I}^{\prime}$ with additive and identically ordered valuations, such that if $\mathcal{I}^{\prime}$ admits an MMS allocation $\mathcal{A}^{\prime}$ then, in fact, $\mathcal{I}$ also admits an MMS allocation $\mathcal{A}$ [^7] See [6] for an approximation-preserving version of this result. The reduction works by deriving from $\mathcal{A}^{\prime}$ a picking sequence over the agents and then showing that greedily assigning goods in $\mathcal{I}$, via this sequence, leads to an MMS allocation $\mathcal{A}$.
At the same time, we note that Theorem 3.4 generalizes to additive valuations if one were to relinquish the requirement that each agent $i$’s bundle, $A_{i}$, must be a subset of $[m]$. That is, in contrast to the rest of the paper, if $A_{i}$ is allowed to be a multiset, in which each extra copy of any good $g$ fetches an additional value of $v_{i}(g)$, then the reduction from [7] applies and the guarantee given in Theorem 3.4 extends to all additive valuations.
3.3 Additive Valuations
Theorem 3.9.
Proof.